Solve Problems Using Linear Systems
17. Ian flew his airplane at best cruise speed for 2 h, then at economy cruise speed for 3 h, covering a total of 850 km. On the following day, he flew a tbest cruise speed for 3 h and at economy cruise speed for 2 h, covering a total of 900 km. Find the best cruise speed and the economy cruise speed for Ian's airplane.
(The answer is: best cruise speed 200 km/h, economy cruise speed 150 km/h, but i dont know how to solve this problem?) Thank you very much!!!

Answer by scott8148(6628)

"Ian flew his airplane at best cruise speed for 2 h, then at economy cruise speed for 3 h, covering a total of 850 km"
___ 2b+3c=850

"he flew a tbest cruise speed for 3 h and at economy cruise speed for 2 h, covering a total of 900 km"
___ 3b+2c=900

multiplying the 1st equation by -2 and the 2nd equation by 3 will allow you to ELIMINATE the c variable by addition

-2(2b+3c=850) ___ -4b-6c=-1700
3(3b+2c=900) ___ 9b+6c=2700
adding equations ___ 5b=1000 ___ b=200

substituting ___ 2(200)+3b=850 ___ 3b=450 ___ b=150